# ParsNets: A Parsimonious Orthogonal and Low-Rank Linear Networks for Zero-Shot Learning

### 15 Dec 2023 v1

@@ 1 Introduction
Figure 1: Subspace visualization of AWA2 (Xian et al., 2018a) on three random classes.
@@
In ZSL/GZSL, since the trained model has no observation of any sample from unseen classes, the mapping function is inherently biased towards seen classes (Fu et al., 2015; Guo et al., 2023b), i.e., the mapped features are usually overfitted to clusters near seen classes, hence can hardly infer to unseen classes with satisfactory performance. To relieve the domain-biased overfitting, especially for GZSL, existing methods usually resorted to learning more representative features having less gap between seen/unseen classes. Some widely used approaches include 1) visual-semantics alignment, for smooth knowledge transfer (Schonfeld et al., 2019; Guo & Guo, 2020); 2) generative methods (i.e., synthesize mimic samples of unseen classes), for training a holistic model (Huang et al., 2019; Zhao et al., 2022); and 3) fine-grained methods, for extracting more generalizable features (Huynh & Elhamifar, 2020; Guo et al., 2023b).
@@
Moreover, it can be noted that existing methods mostly rely on a series of complex deep models to extract and fuse comprehensive features for superior recognition (Xie et al., 2019; Huynh & Elhamifar, 2020; Wang et al., 2018; Guo et al., 2023b). As a result, such training and deployment can be costly in terms of both computing and memory overhead. In this regard, we make an assumption that ZSL/GZSL can be more favorable to a scenario associated with resource-constrained devices due to the low or even zero data requirements. Such a scenario can also well align with ubiquitous devices and data in real-world applications. However, as far as we know, nearly no research has investigated lightweight ZSL/GZSL models deployed on resource-constrained devices.
In this paper, we suggest that the above generalization and lightweight requirements can be jointly achieved by our properly designed parsimonious-yet-efficient network refactoring framework, namely, ParsNets. Specifically, we utilize a set of base linear networks to estimate the nonlinear visual-semantics mapping function that usually involves complex deep models, wherein, each base linear network can correspond to different components of the semantic space shared by both seen and unseen classes (Figure LABEL:fig1). To encourage the learned features to be most discriminative from each other and generalizable to novel concepts, we enforce a low-rank structure to the features of data samples from the same class and a high-rank structure to the features of data samples from all different classes. Hence, intra-class samples can reside in the same linear subspace, and meanwhile, inter-class subspaces can be orthogonal from each other. Such constructed maximal margin geometry is expected to facilitate a smooth knowledge transfer between seen and unseen classes since no entanglement exists. Moreover, to further encourage the model’s adaptability and counterbalance over/under-fittings, we employ a set of sample-wise indicators to select a sparse subset of these base linear networks to form a composite predictor for each sample, thus the global nonlinearity can be collapsed into sparse local linearities to further reduce the computing complexity.
@@ 2.2 Domain-Bias Problem
In recent years, extensive efforts have been made to relieve the domain bias caused by the disjoint train (seen) and test (unseen) classes. For example, some methods tried to align the visual and semantic features within the visual-semantics mapping function to enable a smooth knowledge transfer (Zhang & Saligrama, 2015; Schonfeld et al., 2019; Guo & Guo, 2020). Differently, some other methods proposed to synthesize mimical samples conditioned on unseen class semantic descriptors and jointly train a holistic model combined with seen class samples (Huang et al., 2019; Zhao et al., 2022; Feng et al., 2022). In contrast, some recent methods focused on fine-grained elements or key points within samples to extract more generalizable features between seen/unseen classes (Xie et al., 2019; Huynh & Elhamifar, 2020; Guo et al., 2023b). Despite the relief from the domain-biased overfitting, we observe that underfitting can arise and fluctuate the training towards mediocre convergence for both seen/unseen classes. [This paper]: We construct the maximal margin geometry along with the network refactoring to encourage the discrimination of learned features, and utilize sample-wise indicators to enable a sparse composition of base linear networks. Such regularization can facilitate the generalization towards unseen classes, and explore a better balance between over/under-fittings.
@@ 3.4 Orthogonality and Low-Rankness
PSVMA (Liu et al., 2023a) CVPR ${}^{\prime}{23}$ 79.4 72.9 66.5 45.9
GKU (Guo et al., 2023b) AAAI ${}^{\prime}{23}$ 76.9 - -
@@
PSVMA (Liu et al., 2023a) CVPR ${}^{\prime}{23}$ 73.6 77.3 75.4 70.1 77.8 73.8 61.7 45.3 52.3 - - -
GKU (Guo et al., 2023b) AAAI ${}^{\prime}{23}$ 52.3 71.1 60.3 - - - - - -
@@ 4.2 Comparison of ZSL Performance
We compare the proposed ParsNets with 16 state-of-the-art deep learning-based competitors in the classic ZSL scenario and report the multi-way classification accuracy in Table 1. It can be observed from the results that our method outperforms most deep learning-based competitors on all datasets. For example, DGZ (Chen et al., 2023) and TDCSS (Feng et al., 2022) are the most two powerful generative model-based ZSL competitors that can achieve 80.1% and 61.1% recognition accuracy on CUB-200. In contrast, despite as a non-deep method, our method obtains 0.2% and 19.2% higher performance with much lower computing cost. On the other hand, as a complex graph fine-grained method that utilized powerful GNNs as the sample representation, GKU (Guo et al., 2023b) achieves 76.9% on CUB-200. In contrast, our method obtains a much better performance of 80.3% with/on only the resource-constrained device.
@@ 4.3 Comparison of GZSL Performance
As shown in Table 2, we compare the proposed ParsNets with 15 state-of-the-art deep learning-based competitors in the GZSL scenario and report the average per-class prediction accuracy on unseen classes (U), seen classes (S), and their Harmonic Mean (H). We can observe that our method also constantly outperforms other deep learning-based competitors by even improved margins than that of the ZSL scenario on all datasets. Specifically, in GZSL, most competitors can hardly achieve balanced performance in both seen and unseen classes due to the domain biased over/under-fittings problem. For example, in Zhu et al. (Zhu et al., 2019) and APNet (Liu et al., 2020), there exist 49.5% and 42.0% margins between the accuracy of seen and unseen classes in AWA2 and aPY, respectively, thus the overall performance, i.e, Harmonic Mean, is significantly poor for real-world application. Moreover, even some of the most powerful competitors such as PSVMA (Liu et al., 2023a) and DGZ (Chen et al., 2023) can still have a nonnegligible margin, i.e., 16.4% and 12.3% in SUN and AWA2, respectively. In contrast, due to the utilization of the proposed sample-wise composite semantic predictor and the constructed maximal margin geometry, our method can significantly relieve the domain biased over/under-fittings problem and obtains a more balanced performance in both seen and unseen classes.

### 21 Dec 2023 v2

@@ Abstract
Figure 1: Subspace visualization of AWA2 (Xian et al., 2018a) on three random classes. The orthogonality indicates the maximum separability.
1 Introduction
@@
In ZSL/GZSL, since the trained model has no observation of any sample from unseen classes, the mapping function is inherently biased towards seen classes (Fu et al., 2015; Guo et al., 2023), i.e., the mapped features are usually overfitted to clusters near seen classes, hence can hardly infer to unseen classes with satisfactory performance. To relieve the domain-biased overfitting, especially for GZSL, existing methods usually resorted to learning more representative features having less gap between seen/unseen classes. Some widely used approaches include 1) visual-semantics alignment, for smooth knowledge transfer (Schonfeld et al., 2019; Guo & Guo, 2020); 2) generative methods (i.e., synthesize mimic samples of unseen classes), for training a holistic model (Huang et al., 2019; Zhao et al., 2022); and 3) fine-grained methods, for extracting more generalizable features (Huynh & Elhamifar, 2020; Guo et al., 2023).
@@
Moreover, it can be noted that existing methods mostly rely on a series of complex deep models to extract and fuse comprehensive features for superior recognition (Xie et al., 2019; Huynh & Elhamifar, 2020; Wang et al., 2018; Guo et al., 2023). As a result, such training and deployment can be costly in terms of both computing and memory overhead. In this regard, we make an assumption that ZSL/GZSL can be more favorable to a scenario associated with resource-constrained devices due to the low or even zero data requirements. Such a scenario can also well align with ubiquitous devices and data in real-world applications. However, as far as we know, nearly no research has investigated lightweight ZSL/GZSL models deployed on resource-constrained devices.
In this paper, we suggest that the above generalization and lightweight requirements can be jointly achieved by our properly designed parsimonious-yet-efficient network refactoring framework, namely, ParsNets. Specifically, we utilize a set of base linear networks to estimate the nonlinear visual-semantics mapping function that usually involves complex deep models, wherein, each base linear network can correspond to different components of the semantic space shared by both seen and unseen classes. To encourage the learned features to be most discriminative from each other and generalizable to novel concepts, we enforce a low-rank structure to the features of data samples from the same class and a high-rank structure to the features of data samples from all different classes. Hence, intra-class samples can reside in the same linear subspace, and meanwhile, inter-class subspaces can be orthogonal from each other. Such constructed maximal margin geometry is expected to facilitate a smooth knowledge transfer between seen and unseen classes since no entanglement exists. Moreover, to further encourage the model’s adaptability and counterbalance over/under-fittings, we employ a set of sample-wise indicators to select a sparse subset of these base linear networks to form a composite predictor for each sample, thus the global nonlinearity can be collapsed into sparse local linearities to further reduce the computing complexity.
@@ 2.2 Domain-Bias Problem
In recent years, extensive efforts have been made to relieve the domain bias caused by the disjoint train (seen) and test (unseen) classes. For example, some methods tried to align the visual and semantic features within the visual-semantics mapping function to enable a smooth knowledge transfer (Zhang & Saligrama, 2015; Schonfeld et al., 2019; Guo & Guo, 2020). Differently, some other methods proposed to synthesize mimical samples conditioned on unseen class semantic descriptors and jointly train a holistic model combined with seen class samples (Huang et al., 2019; Zhao et al., 2022; Feng et al., 2022). In contrast, some recent methods focused on fine-grained elements or key points within samples to extract more generalizable features between seen/unseen classes (Xie et al., 2019; Huynh & Elhamifar, 2020; Guo et al., 2023). Despite the relief from the domain-biased overfitting, we observe that underfitting can arise and fluctuate the training towards mediocre convergence for both seen/unseen classes. [This paper]: We construct the maximal margin geometry along with the network refactoring to encourage the discrimination of learned features, and utilize sample-wise indicators to enable a sparse composition of base linear networks. Such regularization can facilitate the generalization towards unseen classes, and explore a better balance between over/under-fittings.
@@ 3.4 Orthogonality and Low-Rankness
PSVMA (Liu et al., 2023) CVPR ${}^{\prime}{23}$ 79.4 72.9 66.5 45.9
GKU (Guo et al., 2023) AAAI ${}^{\prime}{23}$ 76.9 - -
@@
PSVMA (Liu et al., 2023) CVPR ${}^{\prime}{23}$ 73.6 77.3 75.4 70.1 77.8 73.8 61.7 45.3 52.3 - - -
GKU (Guo et al., 2023) AAAI ${}^{\prime}{23}$ 52.3 71.1 60.3 - - - - - -
@@ 4.2 Comparison of ZSL Performance
We compare the proposed ParsNets with 16 state-of-the-art deep learning-based competitors in the classic ZSL scenario and report the multi-way classification accuracy in Table 1. It can be observed from the results that our method outperforms most deep learning-based competitors on all datasets. For example, DGZ (Chen et al., 2023) and TDCSS (Feng et al., 2022) are the most two powerful generative model-based ZSL competitors that can achieve 80.1% and 61.1% recognition accuracy on CUB-200. In contrast, despite as a non-deep method, our method obtains 0.2% and 19.2% higher performance with much lower computing cost. On the other hand, as a complex graph fine-grained method that utilized powerful GNNs as the sample representation, GKU (Guo et al., 2023) achieves 76.9% on CUB-200. In contrast, our method obtains a much better performance of 80.3% with/on only the resource-constrained device.
@@ 4.3 Comparison of GZSL Performance
As shown in Table 2, we compare the proposed ParsNets with 15 state-of-the-art deep learning-based competitors in the GZSL scenario and report the average per-class prediction accuracy on unseen classes (U), seen classes (S), and their Harmonic Mean (H). We can observe that our method also constantly outperforms other deep learning-based competitors by even improved margins than that of the ZSL scenario on all datasets. Specifically, in GZSL, most competitors can hardly achieve balanced performance in both seen and unseen classes due to the domain biased over/under-fittings problem. For example, in Zhu et al. (Zhu et al., 2019) and APNet (Liu et al., 2020), there exist 49.5% and 42.0% margins between the accuracy of seen and unseen classes in AWA2 and aPY, respectively, thus the overall performance, i.e, Harmonic Mean, is significantly poor for real-world application. Moreover, even some of the most powerful competitors such as PSVMA (Liu et al., 2023) and DGZ (Chen et al., 2023) can still have a nonnegligible margin, i.e., 16.4% and 12.3% in SUN and AWA2, respectively. In contrast, due to the utilization of the proposed sample-wise composite semantic predictor and the constructed maximal margin geometry, our method can significantly relieve the domain biased over/under-fittings problem and obtains a more balanced performance in both seen and unseen classes.